OpenStudy (anonymous):

what is: (143(5) + 67)5 ___________ A. 1950 B. 1955 C. 1956

OpenStudy (mervatm):

143*5=715 715+67=782 782*5=3910

OpenStudy (pina_colada123):

thats what got 2

OpenStudy (cecil_the_weasel):

I agree with MervatM. The answer is 3910.

OpenStudy (rocker4744):

same